If N= 2^4*3^5*5^2*7^3Then find number of factors of the form 4n+2.Also find number of factors of the form 8n+2

Grade:11

1 Answers

Meet
6 years ago
4n+2=2(2n+1), it means u can write N as 2(3^0+3^1+.......+3^5)(5^0+.....+5^2)(7^0+..+7^3)= 1×6×3×4= 72.